OpenStudy (anonymous):

convert 110 degrees to radians cant remember which one is 180/pie or pie/180

OpenStudy (anonymous):

multiply by pi, divide by 180 second one

OpenStudy (anonymous):

or even pie! mmmmm pie...

OpenStudy (anonymous):

in this case you get \[\frac{11\pi}{18}\]

OpenStudy (anonymous):

multiply with \[\pi\] and divide with 180